Where is zero field occur between two opposite charges of sme magnitude
1
Expert's answer
2021-09-09T10:49:33-0400

"E_1+E_2=E=0,"

"E_1=-E_2,"

"q_1>0,~q_2<0,"

"\frac{kq_1}{x^2}=-\frac{kq_2}{(r-x)^2},"

"\sqrt{-\frac{q_1}{q_2}}=\frac x{r-x}," "\implies"

"x=\frac{r\sqrt{-\frac{q_1}{q_2}}}{1+\sqrt{-\frac{q_1}{q_2}}}=\frac{r\sqrt{q_1}}{\sqrt{q_1}+\sqrt{-q_2}}."